# A Diffused Background from Axion-like Particles in the Microwave Sky

### v1

@@ Abstract
The nature of dark matter is an unsolved cosmological problem and axions are one of the weakly interacting cold dark matter candidates. Axions or ALPs (Axion-like particles) are pseudo-scalar bosons predicted by beyond-standard model theories. The weak coupling of ALPs with photons leads to the conversion of CMB photons to ALPs in the presence of a transverse magnetic field. If they have the same mass as the effective mass of a photon in a plasma, the resonant conversion would cause a polarized spectral distortion leading to temperature fluctuations with the distortion spectrum. The probability of resonant conversion depends on the properties of the cluster such as the magnetic field, electron density, and its redshift. We show that this kind of conversion can happen in numerous unresolved galaxy clusters up to high redshifts, which will lead to a diffused polarised anisotropy signal in the microwave sky. The spectrum of the signal and its shape in the angular scale will be different from the lensed CMB polarization signal. This new polarised distortion spectrum will be correlated with the distribution of clusters in the universe and hence, with the large-scale structure. The spectrum can then be probed using its spectral and spatial variation with respect to the CMB and various foregrounds. An SNR of $\sim$ 4.36 and $\sim$ 93.87 are possible in the CMB-S4 145 GHz band and CMB-HD 150 GHz band respectively for a photon-ALPs coupling strength of $\mathrm{g_{a\gamma}=10^{-12}\,GeV^{-1}}$ using galaxy clusters beyond redshift z $=1$. The same signal would lead to additional RMS fluctuations of $\sim\mathrm{7.5\times 10^{-2}\,\mu K}$ at 145 GHz. In the absence of any signal, future CMB experiments such as Simons Observatory (SO), CMB-S4, and CMB-HD can put constraints on the coupling strength better than current bounds from particle physics experiment CERN Axion Solar Telescope (CAST).
@@ 1 Introduction
The Cosmic Microwave Background (CMB) is the primordial radiation that surrounds us and is a remnant of the hot Big Bang in the early universe. The initially tightly coupled radiation and baryons were decoupled due to the expansion of the universe. When the optical depth at the time of decoupling was significantly lowered, the photons were able to travel large distances and left the photon-baryon fluid at the last scattering surface at a redshift z $\approx 1089$ [1]. These photons are now observed as the CMB and are highly redshifted with a monopole temperature of 2.7255 K [2]. Many processes affect the CMB at higher multipoles like Doppler shift, lensing, scattering, etc. The power spectrum of the CMB is very well known from the correlation between temperature and polarization power spectra [3]. Cosmic Background Explorer (COBE), Wilkinson Microwave Anisotropy Probe (WMAP) and Planck have provided invaluable information about the CMB and our universe. Several upcoming experiments such as Simons Observatory (SO)[4], CMB-S4[5], CMB-HD [6] will be making even higher resolution CMB observations in the coming years, with an emphasis on probing the polarized CMB and spectral distortions, which refer to the departures from Planck black-body spectrum [7].
The CMB photons as they pass through galaxy clusters, can undergo conversions to axion-like particles (ALPs) if ALPs exist in the universe, irrespective of whether they constitute a fraction of dark matter. These conversions can be resonant or non-resonant and are well studied [8, 9, 10, 11]. The resonant conversions are dominant and require the effective masses of photon and ALP to be equal. This sets the stage for the probability of conversion that depends on the magnetic field and electron density profile, as well as the redshift of the cluster. The conversion leads to a new type of polarized spectral distortion in the CMB. With the ongoing and upcoming experiments, we will be able to gain further insight into the anisotropies and distortions in the CMB. In this paper, we study the capabilities of these experiments in being able to probe this ALP distortion signal from unresolved galaxy clusters.
@@
The motivation behind searching for the diffused ALP signal is highlighted in Sec. 2, followed by the CMB photon to ALP resonant conversion in galaxy clusters in Sec. 3. The ALP background, along with achievable SNR in various experiments after cleaning are analyzed in 4. The estimator for the ALP power spectrum and the related covariance is mentioned in Sec. 5, followed by a spectral and comparison of ALP diffused spectrum with CMB and foregrounds in Sec. 6 . The constraints obtained on ALP coupling constant using ILC are mentioned in Sec. 8. Sec. 9 summarizes the need and the techniques that can be used to increase the detectability of this faint background signal. The power spectrum of the ALP signal produced in a single cluster is studied in Appendix A. The variation of ALP spectrum due to contribution for very high redshifts is explained in Appendix B. The derivation for the map-based power spectrum estimator and the bounds on coupling constant using the template matching of foregrounds is provided in Appendix C and D respectively. We have used natural units ($\mathrm{\hbar=1,c=1,k_{B}=1}$) in most places, until explicitly mentioned. We have used the cosmological parameters from Planck 2015 results [20].
@@ 2 Motivation
To obtain bounds on the background spectrum, a comparison between the observed power spectrum of the microwave sky and that of the fiducial (no ALPs case) power spectrum is required. The background spectrum will contribute to the residual of the two spectrums. Using the covariance for the observed sky, one can obtain bounds on the diffused ALP spectrum from unresolved clusters.
@@
With the upcoming high-resolution CMB experiments such as the SO, CMB-S4, and the proposed CMB-HD, we will be able to estimate the ALP power spectrum, based on its frequency and spatial dependence, which is very distinct as compared to the spectra from other phenomena, to obtain constraints on the photon-ALP coupling constant $\mathrm{g_{a\gamma}}$.
@@ 3 CMB photon - ALP resonant conversion in Galaxy clusters
The CMB black-body radiation peaks at 160.2 GHz, with very tiny fluctuations. Also, it is anisotropic with spatial fluctuations in temperature of the order of $10^{-5}$. Only about $\sim 5\%$ of the CMB photons are linearly polarized. Also, there are spectral distortions in the CMB, which represent its deviations from a black-body spectrum, due to the absorption or emission of photons at different frequencies. These arise due to phenomena like the $\mu$ and $y$ distortions. Earlier the photon-baryon fluid was efficiently thermalized via processes like the Compton, double Compton and bremsstrahlung [21]. Energy release via particle-decays and primordial black hole evaporation gets redistributed. When these processes that redistribute the energy and photons start becoming inefficient due to the Hubble expansion, distortions start setting in, which can be probed to study the early universe physics [22]. The CMB polarization spectrum is mainly affected by Thomson scattering and gravitational lensing, and also exhibits an independence from the frequency of observation [23].
@@
$\mathrm{\mathcal{L}_{int}=-\frac{g_{a\gamma}F_{\mu\nu}\tilde{F}^{\mu\nu}a}{4}=g_{a\gamma}E\cdot B_{ext}a.}$ (3.1)
@@
$\mathrm{m_{a}=m_{\gamma}=\frac{\hbar\omega_{p}}{c^{2}}\approx\frac{\hbar}{c^{2}}\sqrt{n_{e}e^{2}/m_{e}\epsilon_{0}},}$ (3.2)
here $\mathrm{\omega_{p}}$ is the plasma frequency and $\mathrm{n_{e}}$ is the electron density at the location.
So, the electron density at a location in the cluster determines the ALP mass that can be formed at that location. For a spherically symmetric electron density profile, ALPs of a particular mass will be formed in a spherical shell in the cluster. This shell will be projected as a disk in 2D. Higher mass ALPs shall be formed near the center of the cluster with a disk of lower angular size as compared to the lower mass ALPs. The ALPs being probed are in the mass range $10^{-15}-10^{-11}$ eV, which depends on the electron density of the cluster. The dispersion relation for photon-ALP conversion is given as (here $\mathrm{B_{t}}$ refers to the transverse magnetic field):
$\mathrm{2\omega(\omega-k)=-\omega(\Delta_{e}+\Delta_{a})\pm\omega\Delta_{osc}=\frac{m_{a}^{2}+m_{\gamma}^{2}}{2}\pm\left[\left(\frac{m_{a}^{2}-m_{\gamma}^{2}}{2}\right)^{2}+\omega^{2}g_{a\gamma}^{2}B_{t}^{2}\right]^{1/2}},$ (3.3)
@@
$\mathrm{\Delta_{e}\approx\omega_{p}^{2}/2\omega},$
$\mathrm{\Delta_{a}=-m_{a}^{2}/2\omega},$
$\mathrm{\Delta_{a\gamma}=g_{a\gamma}B_{t}/2},$
$\mathrm{\Delta_{osc}^{2}=(\Delta_{a}-\Delta_{e})^{2}+4\Delta_{a\gamma}^{2}}.$
@@
$\mathrm{\gamma_{ad}=\frac{\Delta_{osc}^{2}}{|\nabla\Delta_{e}|}=\left|\frac{2g_{a\gamma}^{2}B_{t}^{2}\nu(1+z)}{\nabla\omega_{p}^{2}}\right|}.$ (3.4)
If the resonant condition is satisfied ($\mathrm{m_{a}=m_{\gamma}}$), in the adiabatic case ($\mathrm{\gamma_{ad}}>>1$), the photon-ALP conversion will definitely occur. In the non-adiabatic case ($\mathrm{\gamma_{ad}}<<1$), the conversion will be probabilistic with the probability given as $\mathrm{1-p}$, where $\mathrm{p}$ refers to the probability of branch shift when passing from a high density to low density region or vice-versa. This is given as [9]:
$\mathrm{p=e^{-\pi\gamma_{ad}/2}}.$ (3.5)
@@
$\mathrm{P(\gamma\rightarrow a)=2p(1-p)=2e^{-\pi\gamma_{ad}/2}(1-e^{-\pi\gamma_{ad}/2}),}$ (3.6)
where $\mathrm{1-p}$ is the probability of conversion at the resonant location.
In the non-adiabatic limit ($\mathrm{\gamma_{ad}}<<1$), this is approximated as:
$\mathrm{P(\gamma\rightarrow a)\approx\pi\gamma_{ad}}.$
The change in intensity due to this conversion in the CMB is given as:
$\mathrm{\Delta I_{\nu}^{\alpha}=P(\gamma\rightarrow a)I_{cmb}(\nu)\approx\pi\gamma_{ad}\left(\frac{2h\nu^{3}}{c^{2}}\right)\frac{1}{e^{h\nu/kT_{cmb}}-1}.}$ (3.7)
@@ 4.1 Halo modeling of matter in the universe
$\mathrm{\rho_{m}(x)=[1+\delta_{m}(x)]\bar{\rho_{m}}},$ (4.1)
with $\mathrm{\bar{\rho_{m}}}$ being the mean matter density. The dark matter is the dominant constituent of the matter density in the universe. The distribution of these halos can be assumed to follow the matter distribution in the universe. This gives the matter density field from a superposition of halos $\mathrm{{}^{\prime}i^{\prime}}$ of masses $\mathrm{M_{i}}$( [1]):
$\mathrm{\rho_{m}(x)=\sum_{i}\rho_{h}(|x-x_{i}|,M_{i})},$ (4.2)
with $\mathrm{\rho_{h}}$ being the density profile of the halo. These halos are assumed to only be interacting gravitationally and hence, their properties only depend on their masses. These are assumed to have undergone spherical collapse and subsequent virialization. Their masses are typically defined as the mass within the radius at which the density of the halo becomes about 200 times the critical density of the universe $\mathrm{\rho_{c}}$. The profiles start steepening beyond this radius.
@@
$\mathrm{\delta_{m}(x)=\int d\,ln\,M\frac{M}{\rho_{m}}\frac{dn}{d\,ln\,M}\int d^{3}x^{\prime}\delta_{h}(x^{\prime},M)y(|x-x^{\prime}|,M)},$ (4.3)
where $\mathrm{\delta_{h}}$ accounts for the variations in the mass function, i.e.,
$\mathrm{\frac{dn(x)}{dM}=[1+\delta_{h}(x,M)]\frac{dn}{dM}}.$ (4.4)
@@
$\mathrm{dn/dM=f(\sigma)\frac{\rho_{c}\Omega_{m}}{M}\frac{d\,ln\sigma^{-1}}{dM}}.$ (4.5)
Here $\mathrm{\Omega_{m}}$ is the matter density parameter. The $\sigma$ represents the RMS deviation in the initial density fluctuation field smoothed with a tophat filter, and $\mathrm{f(\sigma)}$ is the halo multiplicity function. The Tinker mass function [28] uses the multiplicity function with four free parameters ($\mathrm{d=1.97,e=1.00,f=0.51,g=1.228}$) and a normalization ($\mathrm{B}=0.482$) and the values are set to those for $\mathrm{M_{200m}}$. The multiplicity function reads
$\mathrm{f(\sigma)=B\left[\left(\frac{\sigma}{e}\right)^{-d}+\sigma^{-f}\right]exp(-g/\sigma^{2}).}$ (4.6)
The halo mass function decreases with increasing redshifts and increasing masses, which leads to more heavier clusters at lower redshifts. This is reflected in the probability of clusters of higher masses being low as compared to those of lower mass in the redshift range of 0.5 to 7. Also, the slope becomes steeper with increasing redshift at higher masses. This indicates that the proportion of high to low-mass clusters is greater at low redshifts as compared to higher ones. In the halo model, the non-linear matter power spectrum can be written as the sum of the contributions from one and two-halo terms, i.e.,
$\mathrm{P(k)=P_{1h}(k)+P_{2h}(k)}.$ (4.7)
@@
$\mathrm{\xi_{mm}(|\textbf{x}_{1}-\textbf{x}_{2}|)=\langle\delta_{m}(\textbf{x}_{1})\delta_{m}(\textbf{x}_{2})\rangle}.$ (4.8)
@@
$\mathrm{\xi_{mm}(r)=\frac{1}{2\pi^{2}}\int dkk^{2}P(k)\frac{sin(kr)}{kr}.}$ (4.9)

### v3

@@ Abstract
The nature of dark matter is an unsolved cosmological problem and axions are one of the weakly interacting cold dark matter candidates. Axions or ALPs (Axion-like particles) are pseudo-scalar bosons predicted by beyond-standard model theories. The weak coupling of ALPs with photons leads to the conversion of CMB photons to ALPs in the presence of a transverse magnetic field. If they have the same mass as the effective mass of a photon in a plasma, the resonant conversion would cause a polarized spectral distortion leading to temperature fluctuations with the distortion spectrum. The probability of resonant conversion depends on the properties of the cluster such as the magnetic field, electron density, and its redshift. We show that this kind of conversion can happen in numerous unresolved galaxy clusters up to high redshifts, which will lead to a diffused polarised anisotropy signal in the microwave sky. The spectrum of the signal and its shape in the angular scale will be different from the lensed CMB polarization signal. This new polarised distortion spectrum will be correlated with the distribution of clusters in the universe and hence, with the large-scale structure. The spectrum can then be probed using its spectral and spatial variation with respect to that of the CMB and various foregrounds. An SNR of $\sim$ 4.36 and $\sim$ 93.87 are possible in the CMB-S4 145 GHz band and CMB-HD 150 GHz band respectively for a photon-ALPs coupling strength of $g_{a\gamma}=10^{-12}\,\mathrm{GeV}^{-1}$ using galaxy clusters beyond redshift $z=1$. The same signal would lead to additional RMS fluctuations of $\sim\mathrm{7.5\times 10^{-2}\,\mu K}$ at 145 GHz. In the absence of any signal, future CMB experiments such as Simons Observatory (SO), CMB-S4, and CMB-HD can put constraints on the coupling strength better than current bounds from particle physics experiment CERN Axion Solar Telescope (CAST).
@@ 1 Introduction
The Cosmic Microwave Background (CMB) is the primordial radiation that surrounds us and is a remnant of the hot Big Bang in the early universe. The initially tightly coupled radiation and baryons were decoupled due to the expansion of the universe. When the optical depth at the time of decoupling was significantly lowered, the photons were able to travel large distances and left the photon-baryon fluid at the last scattering surface at a redshift $z\approx 1089$ [1]. These photons are now observed as the CMB and are highly redshifted with a monopole temperature of 2.7255 K [2]. Many processes affect the CMB at higher multipoles like Doppler shift, lensing, scattering, etc. The power spectrum of the CMB is very well known from the correlation between temperature and polarization power spectra [3]. Cosmic Background Explorer (COBE), Wilkinson Microwave Anisotropy Probe (WMAP) and Planck have provided invaluable information about the CMB and our universe. Several upcoming experiments such as Simons Observatory (SO)[4], CMB-S4[5], CMB-HD [6] will be making even higher resolution CMB observations in the coming years, with an emphasis on probing the polarized CMB and spectral distortions [7, 8, 9, 10], which refer to the departures from Planck black-body spectrum [11, 12].
Axions are mass possessing pseudo Nambu-Goldstone bosons that solve the strong CP problem and possess mass, making them of great interest to solve the dark matter problem [13, 14, 15, 16]. Their production requires high energies in particle colliders [17], hence probing their weak coupling with photons is a great way to detect them. In various studies, the existence of axions or ALPs will lead to spectral distortions in the CMB [18, 19, 20].
The CMB photons as they pass through galaxy clusters, can undergo conversions to ALPs if ALPs exist in the universe, irrespective of whether they constitute a fraction of dark matter [21, 22, 23, 24, 25]. These conversions can be resonant or non-resonant and are well studied [26, 27, 28, 29, 30]. The resonant conversions are dominant and require the effective masses of photon and ALP to be equal. This sets the stage for the probability of conversion that depends on the magnetic field and electron density profile, as well as the redshift of the cluster. The conversion leads to a new type of polarized spectral distortion in the CMB. With the ongoing and upcoming experiments, we will be able to gain further insight into the anisotropies and distortions in the CMB. In this paper, we study the capabilities of these experiments in being able to probe this ALP distortion signal from unresolved galaxy clusters.
@@
The motivation behind searching for the diffused ALP signal is highlighted in Sec. 2, followed by the CMB photon to ALP resonant conversion in galaxy clusters in Sec. 3. The ALP background and its sources of variation are analyzed in Sec. 4. The estimator for the ALP power spectrum and the related covariance is mentioned in Sec. 5, followed by a spectral comparison of ALP diffused spectrum with CMB and foregrounds, and the achievable SNR for various experiments in Sec. 6. The constraints obtained on ALP coupling constant using ILC are mentioned in Sec. 8. Sec. 9 summarizes the need and the techniques that can be used to increase the detectability of this faint background signal. The power spectrum of the ALP signal produced in a single cluster is studied in Appendix A. The variation of ALP spectrum due to contribution for very high redshifts is explained in Appendix B. The derivation for the map-based power spectrum estimator and the bounds on coupling constant using the template matching of foregrounds is provided in Appendix C and D respectively. We have used natural units ($\hbar=1,c=1,k_{B}=1$) in most places, until explicitly mentioned. We have used the cosmological parameters from Planck 2015 results [39].
@@ 2 Motivation
To obtain bounds on the background spectrum, a comparison between the observed power spectrum of the microwave sky and that of the fiducial (no ALPs case) power spectrum is required. The background spectrum will contribute to the residual of the two spectra. Using the covariance for the observed sky, one can obtain bounds on the diffused ALP spectrum from unresolved clusters.
@@
With the upcoming high-resolution CMB experiments such as the SO, CMB-S4, and the proposed CMB-HD, we will be able to estimate the ALP power spectrum, based on its frequency and spatial dependence, which is very distinct as compared to the spectra from other phenomena, to obtain constraints on the photon-ALP coupling constant $g_{a\gamma}$.
@@ 3 CMB photon - ALP resonant conversion in Galaxy clusters
The CMB black-body radiation peaks at 160.2 GHz, with very tiny fluctuations. Also, it is anisotropic with spatial fluctuations in temperature of the order of $10^{-5}$. Only about $\sim 5\%$ of the CMB photons are linearly polarized. Also, there are spectral distortions in the CMB, which represent its deviations from a black-body spectrum, due to the absorption or emission of photons at different frequencies. These arise due to phenomena like the $\mu$ and $y$ distortions. Earlier the photon-baryon fluid was efficiently thermalized via processes like the Compton, double Compton and bremsstrahlung [40, 11]. Energy release via particle-decays and primordial black hole evaporation gets redistributed. When these processes that redistribute the energy and photons, start becoming inefficient due to the Hubble expansion, distortions start setting in, which can be probed to study the early universe physics [41]. The CMB polarization spectrum is mainly affected by Thomson scattering and gravitational lensing, and also exhibits an independence from the frequency of observation [42].
@@
$\mathcal{L}_{\mathrm{int}}=-\frac{g_{a\gamma}F_{\mu\nu}\tilde{F}^{\mu\nu}a}{4}=g_{a\gamma}E\cdot B_{\mathrm{ext}}a.$ (3.1)
@@
$m_{a}=m_{\gamma}=\frac{\hbar\omega_{p}}{c^{2}}\approx\frac{\hbar}{c^{2}}\sqrt{n_{e}e^{2}/m_{e}\epsilon_{0}},$ (3.2)
here $\omega_{p}$ is the plasma frequency and $n_{e}$ is the electron density at the location.
So, the electron density at a location in the cluster determines the ALP mass that can be formed at that location. For a spherically symmetric electron density profile, ALPs of a particular mass will be formed in a spherical shell in the cluster. This shell will be projected as a disk in 2D. Higher mass ALPs shall be formed near the center of the cluster with a disk of lower angular size as compared to the lower mass ALPs. The ALPs being probed are in the mass range $10^{-15}-10^{-11}$ eV, which depends on the electron density of the cluster. The dispersion relation for photon-ALP conversion is given as (here $B_{t}$ refers to the transverse magnetic field):
$2\omega(\omega-k)=-\omega(\Delta_{e}+\Delta_{a})\pm\omega\Delta_{\mathrm{osc}}=\frac{m_{a}^{2}+m_{\gamma}^{2}}{2}\pm\left[\left(\frac{m_{a}^{2}-m_{\gamma}^{2}}{2}\right)^{2}+\omega^{2}g_{a\gamma}^{2}B_{t}^{2}\right]^{1/2},$ (3.3)
@@
$\begin{split}\Delta_{a}&=-m_{a}^{2}/2\omega\,,\qquad\Delta_{e}\approx\omega_{p}^{2}/2\omega\,,\\ \Delta_{a\gamma}&=g_{a\gamma}B_{t}/2\,,\qquad\Delta_{\mathrm{osc}}^{2}=(\Delta_{a}-\Delta_{e})^{2}+4\Delta_{a\gamma}^{2},\end{split}$ (3.4)
@@
$\gamma_{\mathrm{ad}}=\frac{\Delta_{\mathrm{osc}}^{2}}{|\nabla\Delta_{e}|}=\left|\frac{2g_{a\gamma}^{2}B_{t}^{2}\nu(1+z)}{\nabla\omega_{p}^{2}}\right|.$ (3.5)
If the resonant condition is satisfied ($m_{a}=m_{\gamma}$), in the adiabatic case ($\gamma_{\mathrm{ad}}>>1$), the photon-ALP conversion will definitely occur. In the non-adiabatic case ($\gamma_{\mathrm{ad}}<<1$), the conversion will be probabilistic with the probability given as $1-p$, where $p$ refers to the probability of branch shift when passing from a high density to low density region or vice-versa. This is given as [27]:
$p=e^{-\pi\gamma_{\mathrm{ad}}/2}.$ (3.6)
@@
$P(\gamma\rightarrow a)=2p(1-p)=2e^{-\pi\gamma_{\mathrm{ad}}/2}(1-e^{-\pi\gamma_{\mathrm{ad}}/2}),$ (3.7)
where $1-p$ is the probability of conversion at the resonant location.
In the non-adiabatic limit ($\gamma_{\mathrm{ad}}<<1$), this is approximated as: $P(\gamma\rightarrow a)\approx\pi\gamma_{\mathrm{ad}}$. The change in intensity due to this conversion in the CMB is given as:
$\Delta I_{\nu}^{\alpha}=P(\gamma\rightarrow a)I_{\mathrm{cmb}}(\nu)\approx\pi\gamma_{\mathrm{ad}}\left(\frac{2h\nu^{3}}{c^{2}}\right)\frac{1}{e^{h\nu/k_{B}T_{\mathrm{cmb}}}-1}.$ (3.8)
@@ 4.1 Halo modeling of matter in the universe
$\rho_{m}(x)=[1+\delta_{m}(x)]\bar{\rho_{m}},$ (4.1)
with $\bar{\rho_{m}}$ being the mean matter density. The dark matter is the dominant constituent of the matter density in the universe. The distribution of these halos can be assumed to follow the matter distribution in the universe. This gives the matter density field from a superposition of halos ${}^{\prime}i^{\prime}$ of masses $M_{i}$( [1]):
$\rho_{m}(x)=\sum_{i}\rho_{h}(|x-x_{i}|,M_{i}),$ (4.2)
with $\rho_{h}$ being the density profile of the halo. These halos are assumed to only be interacting gravitationally and hence, their properties only depend on their masses. These are assumed to have undergone spherical collapse and subsequent virialization. Their masses are typically defined as the mass within the radius at which the density of the halo becomes about 200 times the critical density of the universe $\rho_{c}$. The profiles start steepening beyond this radius.
@@
$\delta_{m}(x)=\int\mathrm{d}\,\ln\,M\frac{M}{\rho_{m}}\frac{\mathrm{d}n}{\mathrm{d}\,\ln\,M}\int\mathrm{d}^{3}x^{\prime}\delta_{h}(x^{\prime},M)y(|x-x^{\prime}|,M),$ (4.3)
where $y(x,M)=\rho_{h}(x,M)/M$ is the normalized profile and $\delta_{h}$ accounts for the variations in the mass function, i.e.,
$\frac{\mathrm{d}n(x)}{\mathrm{d}M}=[1+\delta_{h}(x,M)]\frac{\mathrm{d}n}{\mathrm{d}M}.$ (4.4)
@@
$\mathrm{d}n/\mathrm{d}M=f(\sigma)\frac{\rho_{c}\Omega_{m}}{M}\frac{\mathrm{d}\,\ln\sigma^{-1}}{\mathrm{d}M}.$ (4.5)
Here $\Omega_{m}$ is the matter density parameter. The $\sigma$ represents the RMS deviation in the initial density fluctuation field smoothed with a tophat filter, and $f(\sigma)$ is the halo multiplicity function. The Tinker mass function [47] uses the multiplicity function with four free parameters ($d=1.97,e=1.00,f=0.51,g=1.228$) and a normalization ($B=0.482$) and the values are set to those for $M_{200m}$. The multiplicity function reads
$f(\sigma)=B\left[\left(\frac{\sigma}{e}\right)^{-d}+\sigma^{-f}\right]\exp(-g/\sigma^{2}).$ (4.6)
In the halo model, the non-linear matter power spectrum can be written as the sum of the contributions from one and two-halo terms, i.e.,
$P(k)=P_{1h}(k)+P_{2h}(k).$ (4.7)
@@
$\xi_{mm}(|\textbf{x}_{1}-\textbf{x}_{2}|)=\langle\delta_{m}(\textbf{x}_{1})\delta_{m}(\textbf{x}_{2})\rangle.$ (4.8)
@@
$\xi_{mm}(r)=\frac{1}{2\pi^{2}}\int\mathrm{d}k\,k^{2}P(k)\frac{\sin(kr)}{kr}.$ (4.9)